Use a triple integral to determine the volume of the region bounded by

z

=

x

2

+

y

2

and

z

=

x

2

+

y

2

in 1st octant.



1
Expert's answer
2021-07-23T09:54:24-0400

By equating the coordinates "z," we get the following equation:


"\sqrt{x^2+y^2}=x^2+y^2"

"x^2+y^2=0\ or\ x^2+y^2=1"

Using the cylindrical coordinates

"V=\int\int\int_EdV="

"=\displaystyle\int_{0}^{\pi/2}\displaystyle\int_{0}^{1}\displaystyle\int_{r^2}^{r}rdzdrd\theta"

"=\displaystyle\int_{0}^{\pi/2}\displaystyle\int_{0}^{1}(r^2-r^3)drd\theta"

"=\displaystyle\int_{0}^{\pi/2}\bigg[\dfrac{r^3}{3}-\dfrac{r^4}{4}\bigg]\begin{matrix}
 1 \\
 0
\end{matrix}d\theta"

"=\dfrac{1}{12}\big[\theta\big]\begin{matrix}
 \pi/2 \\
 0
\end{matrix}=\dfrac{\pi}{24}(units^3)"

"V=\dfrac{\pi}{24}" cubic units